# Fe contribution to the magnetic anisotropy of $L{\rm 1_{0}}$-ordered FePt thin films studied by angle-dependent x-ray magnetic circular dichroism

### v1

@@ Acknowledgements.
We would like to thank Kenta Amemiya and Masako Suzuki-Sakamaki for valuable technical support at KEK-PF. We would also like to thank Iwao Matsuda for valuable advice and enlightening discussion. This work was supported by Grants-in-Aid for Scientific Research from JSPS (grant Nos. 15H02109, 15K17696, 20K14416, and 22K03535), MEXT Initiative to Establish Next-generation Novel Integrated Circuits Centers (X-NICS) Grant Number JPJ011438, and the National Science and Technology Council of Taiwan under Grant No. 113-2112-M-007-033. The experiment was performed at BL-16A of KEK-PF with the approval of the Photon Factory Program Advisory Committee (proposal Nos. 2016G066, and 2016S2-005). A. F. is an adjunct member of Center for Spintronics Research Network (CSRN), the University of Tokyo, under Spintronics Research Network of Japan (Spin-RNJ), and acknowledges support from the Yushan Fellow Program under the Ministry of Education of Taiwan. T. S. and K. T. are members of CSRN, Tohoku University, under Spin-RNJ.

### v2

@@ Acknowledgements.
We would like to thank Kenta Amemiya and Masako Suzuki-Sakamaki for valuable technical support at KEK-PF. We would also like to thank Iwao Matsuda for valuable advice and enlightening discussion. This work was supported by Grants-in-Aid for Scientific Research from JSPS (grant Nos. 15H02109, 15K17696, 20K14416, and 22K03535), MEXT Initiative to Establish Next-generation Novel Integrated Circuits Centers (X-NICS) Grant Number JPJ011438, and the National Science and Technology Council of Taiwan under Grant No. 113-2112-M-007-033. The experiment was performed at BL-16A of KEK-PF with the approval of the Photon Factory Program Advisory Committee (proposal Nos. 2016G066, and 2016S2-005). A.F. is an adjunct member of Center for Spintronics Research Network (CSRN), the University of Tokyo, under Spintronics Research Network of Japan (Spin-RNJ), and acknowledges support from the Yushan Fellow Program under the Ministry of Education of Taiwan. T.S., S.S. and K.T. are members of CSRN, Tohoku University, under Spin-RNJ.